OpenStudy (anonymous):

no solution, one solution, or infinitely y=5x-4 y=5x-5

OpenStudy (anonymous):

what do the slopes say?

OpenStudy (anonymous):

Both lines have equal slopes.

OpenStudy (anonymous):

ohh, same slope. better check the intercepts...

OpenStudy (anonymous):

No need to check intercepts.

OpenStudy (anonymous):

what do you know about 2 lines that have the same slope?

OpenStudy (anonymous):

really...

OpenStudy (anonymous):

y = 5x + 3 y = 5x + 3 how many solutions?

OpenStudy (anonymous):

Equal slopes means parallel lines. Which means no solution as they will NEVER intersect. The question's 2 lines are y = 5x - 4 and y = 5x - 5

OpenStudy (anonymous):

i'm doing a similar questions so .. 2 lines that have the same slope have no solutions right ? @Easyaspi314

OpenStudy (anonymous):

negative you have to check the intercepts

OpenStudy (anonymous):

Correct. Unless both equations are identical. lol

OpenStudy (anonymous):

so what is it?

OpenStudy (anonymous):

2x + 3y = 4 4x + 6y = 8 better check the x or y intercepts if they have the same slope

OpenStudy (anonymous):

no solution @yyessica_99

OpenStudy (anonymous):

thank you @Little_Kay

OpenStudy (anonymous):

Np

OpenStudy (anonymous):

same slope and different intercepts => no solution sam slope and same intercepts => infinite solutions

OpenStudy (anonymous):

which one is the slope 5x or -4

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

wow what @pgpilot326

OpenStudy (anonymous):

nope, it's the 5 part of the 5x

OpenStudy (anonymous):

lol yeah that's what I meant @pgpilot326

OpenStudy (anonymous):

ok thats all i wanted to know!! @pgpilot326

OpenStudy (anonymous):

2x + 3y = 4 4x + 6y = 8 These 2 equations are identical equations, identical line...as the second equation is the first equation multiplied by 2. So no need to check intercepts, although for many students that may feel more comfortable. But it isnt necessary if the student can notice that they are the same line.